OpenStudy (anonymous):

Help! I am trying to figure out how to isolate the y when I have the equation with a fraction in it.... x/2 + 3y = 4 your help in explaining this one would be greatly appreciated! KJ

OpenStudy (anonymous):

2 * (x/2 + 3y) = 2 *4 x + 6y = 8 6y = 8 - x y = 4/3 - x/6

OpenStudy (saifoo.khan):

y = -x/6 + 4/3